\medskip

\begin{thebibliography}{8}
	

\tiny
	

\bibitem{Aanderaa}
S.O.~Aanderaa, 
\textit{A proof of Higman's embedding theorem using Britton extensions of groups},
Word Probl., Decision Probl. Burnside Probl. Group Theory, Studies Logic Foundations Math. 71 (1973), 1--18.

\bibitem{Atabekyan Gevorkyan 2022}
V.S.~Atabekyan, G.G.~Gevorkyan,
\textit{Central extensions of $n$-torsion groups,}
J. Contemp. Math. Anal., Armen. Acad. Sci. 57 (2022), 1, 26--34. 
	
\bibitem{Belk Hyde Matucci}
J.~Belk, J.~Hyde, F.~Matucci,
\textit{Embeddings $\mathbb Q$ into a finitely presented group}, 
Bull. Amer. Math. Soc. 59 (2022), 561--567. 
\bibitem{Higman Subgroups of fP groups}
G. Higman, 
{\it Subgroups of finitely presented groups}, Proc. R. Soc. Ser. A (1961), 262, 455--475.

\bibitem{HigmanNeumannNeumann}
G. Higman, B. Neumann, Hanna Neumann,
\textit{Embedding theorems for groups}, 
J. London Math. Soc., 3, 24, (1949), 247--254.

\bibitem{Subnormal embedding theorems}
V.\,H. Mikaelian,
\textit{Subnormal embedding theorems for groups}
J. London Math. Soc., 62.2 (2000), 398--406.



\bibitem{Mikaelian 2002}
V.\,H. Mikaelian,
\textit{Two problems on varieties of groups generated by wreath products of groups}, 
International Journal of Mathematics and Mathematical Sciences, 31 (2002), 2, 65--75.


\bibitem{Mikaelian 2003}
V.\,H. Mikaelian,
\textit{On wreath products of finitely generated abelian groups}, 
Advances in Group Theory, Proc. Internat. Research Bimester dedicated to the memory of Reinhold Baer, (Napoli, Italy, May--June, 2002), Aracne, Roma, 2003, 13--24

\bibitem{Mikaelian 2003b}
V.\,H. Mikaelian,
\textit{Infinitely many not locally soluble SI*-groups}, 
Ricerche di Matematica, Univ. Studi Napoli, Naples, 52 (2003), 1--19.


\bibitem{Mikaelian 2004}
V.\,H. Mikaelian,
\textit{On embedding properties of SD-groups}, 
International Journal of Mathematics and Mathematical Sciences, 2004:2 (2004) 65--76. 


\bibitem{Mikaelian 2005}
V.\,H. Mikaelian,
\textit{On a problem on explicit embeddings of the group $\mathbb Q$},
International Journal of Mathematics and Mathematical Sciences, 2005:13 (2005), 2119--2123.


\bibitem{Mikaelian 2007}
V.\,H. Mikaelian,
\textit{On finitely generated soluble non-Hopfian groups, an application to a problem of Neumann}, 
International Journal of Algebra and Computation, 17 (2007), Nos. 5--6, 1107--1113.



\bibitem{metabelian}
V.\,H. Mikaelian, 
\textit{Metabelian varieties of groups and wreath products of abelian groups}, J. Algebra, 2007 (313), 2, 455--485.


\bibitem{On abelian subgroups of}
V.\,H. Mikaelian, A.Yu.~Olshanskii, 
\textit{On abelian subgroups of finitely generated metabelian groups}, J.~Group Theory, 16 (2013), 695--705.



\bibitem{The Criterion of Shmel'kin}
V.\,H. Mikaelian,
\textit{The Criterion of Shmel'kin and Varieties Generated by Wreath Products of Finite Groups},
Algebra and logic, 56, 2, (2017), 108--115.




\bibitem{Subvariety structures}
V.\,H. Mikaelian,
\textit{Subvariety structures in certain product varieties of groups}, J.~Group Theory, 21 (2018), 5, 865--884.


\bibitem{Embeddings using universal words}
V.\,H. Mikaelian,
\textit{Embeddings using universal words in the free group of rank 2}, Sib. Math.~J., 62 (2021), 1, 123--130.

\bibitem{The Higman operations and  embeddings}
V.\,H. Mikaelian,
\textit{The Higman operations and  embeddings of recursive groups}, 
J. Group Theory, DOI: 10.1515/jgth-2021-0095,
\href{https://arxiv.org/abs/2002.09728}{arXiv:2002.09728}

\bibitem{A modified proof for Higman}
V.\,H.~Mikaelian,
\textit{A modified proof for Higman's embedding theorem},
\href{https://arxiv.org/abs/1908.10153}{arXiv:1908.10153}


\bibitem{The conjugacy problem and Higman embeddings}
A.Yu  Ol'shanskii, M. Sapir,
\textit{The conjugacy problem and Higman embeddings}, 
Mem. Amer. Math. Soc. 170 (2004).


\bibitem{Length and area functions on groups and quasi-isometric Higman embeddings}
A.Yu  Ol'shanskii, M. Sapir,
\textit{Length and area functions on groups and quasi-isometric Higman embeddings}, 
International Journal of Algebra and Computation, 11 (2001), 2, 137--170.











\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb,amscd}
\usepackage[bitstream-charter]{mathdesign}
\DeclareSymbolFont{cmsymbols}{OMS}{cmsy}{m}{n}
\SetSymbolFont{cmsymbols}{bold}{OMS}{cmsy}{b}{n}
\DeclareSymbolFontAlphabet{\mathcal}{cmsymbols}

\usepackage[%hypertex,
unicode=true,
plainpages = false,
pdfpagelabels,
bookmarks=true,
bookmarksnumbered=true,
bookmarksopen=true,
breaklinks=true,
backref=false,
colorlinks=true,
linkcolor = blue,
urlcolor  = blue,
citecolor = red,
anchorcolor = green,
hyperindex = true,
hyperfigures
]{hyperref}
\hypersetup{
	pdftitle={Linear algebra, theory and algorithms},
	pdfauthor={V. Mikaelian},
	pdfsubject={Linear Algebra}
}

\title[Higman operations on integer sequences]{\large Higman operations on integer sequences, and embeddings of recursive groups}

\author[V.\,H. Mikaelian]{V.\,H. Mikaelian\\ \lowercase{ v.mikaelian@gmail.com }}

\begin{document}

\maketitle

{\small\noindent This is an extended version of summary of the talk at the \textit{International Conference on Group Theory 
in honor of Victor Mazurov on the occasion of his 80th birthday}, 
Novosibirsk, 2--8 July 2023. The concise version of this report can be found in the talks and communications band at the Conference site:\\
\href{https://sites.google.com/view/mazurov2023}{\tt \footnotesize https://sites.google.com/view/mazurov2023}}

\bigskip
\noindent
The objective of the current talk is to present our very recent article \cite{The Higman operations and  embeddings} in which we for various types of recursive groups discuss possibilities of their \textit{explicit} embeddings into finitely presented groups.

\section{Higman's embedding theorem}
\noindent
In 1961 Higman proved that \textit{a finitely generated group can be embedded in a finitely presented group if and only if it is recursively presented} \cite{Higman Subgroups of fP groups}.
Higman's work is based on specific recursively enumerable sets of integer sequences which in some sense ``code'' the defining relations of groups. 

The algorithm we suggested in \cite{The Higman operations and  embeddings} allows  to list certain wide classes of groups for which Higman's famous embedding construction can also be constructive and explicit. 

\medskip
As a first step, a recursive group 
$G = \langle  A \mathrel{|} R \rangle = \langle a_1,a_2,\ldots \mathrel{|} r_1, r_2,\ldots \rangle $
with recursively enume\-rable relations $r_1, r_2,\ldots$
%($r_k=r_k(a_{i_1},\ldots,a_{i_k})$)
can be constructively embedded into a 
$2$-generator group
$T=\langle b,c 
\mathrel{|} r'_1, r'_2,\ldots 
\rangle$ 
where the relations $r'_1=r'_1(b,c),\;r'_2=r'_2(b,c),\ldots$ are certain words on just two letters $b,c$,  and they also are recursively enumerable (see  \cite{HigmanNeumannNeumann}
for the original embedding theorem and 
 \cite{Embeddings using universal words} for a method of embedding  that preserves the recursive enumeration).
Then for each $r'_n$,\; $n\!=\!1,2,\ldots$, a unique sequence $f_n$ of integers is compiled so that the set $\{r'_1, r'_2,\ldots \}$ of relations
is ``coded'' by means of the set $\mathcal B=\big\{f_1, f_2,\ldots \big\}$ of such sequences. 
Namely, for a relation: 
$$
r'_i  =r'_i(b,c) = b^{n_0}c^{n_1} \cdots b^{n_{2m}}c^{n_{2m+1}}
$$ 
for some $m=m(i)$, and $n_1,\ldots,n_{2m} \neq 0$ (the cases $n_0=0$, or $n_{2m+1}$ are \textit{not} ruled out) we output:
$$
f_i =(n_0, n_1, \ldots ,n_{2m+1}).
$$
Say, for the commutator word: 
$$r_1(b,c)=[b,c]=b^{-1}c^{-1}bc$$ 
we have the sequence: 
$$
f_1=(-1, -1, 1, 1).
$$

Since the transaction from relations set $R$ to sequences set $\mathcal B$ is done via a just few constructive steps, the set $\mathcal B$ also is recursively enumerable.

\medskip
Further for each $f_i\in \mathcal B$ Higman sets some special elements $b_{f_i}$ and $a_{f_i}$ in the free group $F_3 = \langle a,b,c\rangle$ of rank $3$, and using them defines certain respective subgroup $A_{\mathcal B}$ in $F_3$.

Then the so-called \textit{benign} subgroups are defined. One of the key results of \cite{Higman Subgroups of fP groups} is that \textit{$\mathcal B$ is recursively enumerable if and only if  $A_{\mathcal B}$ is benign in $F_3$}.

Finally ``the Higman Rope Trick'' in \cite{Higman Subgroups of fP groups} uses this benign subgroup $A_{\mathcal B}$ to embed $T$, and thus also the initial group $G$, into a finitely presented group.

\section{The explicit constructions}
\noindent
Our note \cite{The Higman operations and  embeddings} and this talk mainly concern the \textit{explicit construction of the recursive sets $\mathcal B$}, so we will discuss them in the sequel. 
The hard part of \cite{Higman Subgroups of fP groups}
is to show that $\mathcal B$ is recursively enumerable if and only if $\mathcal B$ can be constructed by some chain of special operators suggested by Higman:
\begin{equation}
\label{EQ Higman operations}
\iota,\; 
\upsilon,\; 
\rho,\; 
\sigma,\; 
\tau,\; 
\theta,\; 
\zeta,\; 
\pi,\; 
\omega_m.
\end{equation}
And parallel to application of those operations a respective \textit{benign subgroup} is being constructed in the free group $F_3$.

\medskip
In~\cite{Higman Subgroups of fP groups} Higman just relies on \textit{theoretical possibility} for construction of $\mathcal B$ via special operations \eqref{EQ Higman operations}, without any \textit{examples} of such construction for certain particular recursive groups. 
Is it worth noting that after Higman's result there was no attempt to explicitly find constructions of $\mathcal B$ by Higman's operations  for particular groups.
At least, we haven not seen them in the literature.

\medskip
We noticed that such a construction may be a doable task for some classes of groups for which the set $\mathcal B$ obeys certain simple ``combinatorial'' rules.
To explain what we understand under ``combinatorial'' rule let us bring an oversimplified example. Assume $\mathcal B$ consists of sequences $f_i\! =\!(n_0, n_1, \ldots ,n_{2m+1})$ where, say:\\
$n_0$ accepts any value grater then a fixed integer,  \\
$n_1$ accepts any value between two fixed integers, \\
$n_2$ accepts any value less than a fixed integer, \\
$n_3$ accepts arbitrary integer values, \\
$n_4$ and $n_5$ can accept any values but they are opposites of each other, etc. 
(an example will be given for $\mathbb Q$ below).

We noticed that this pattern does occur for many classes of groups, such as, the free abelian, metabelian, soluble, nilpotent groups, the additive group of  rational numbers $\mathbb Q$, the quasicyclic group $\mathbb C_{p^\infty}$, 
divisible abelian groups, etc.

Say, for the group $\mathbb Q$ by Example 3.5 in \cite{The Higman operations and  embeddings} the respective set $\mathcal B$ consists of all tuples $f$ of type:
$$
f_k=
\big(1,-k,-1, -k , -1,\; k,\;  1,\;  k, \; 1,\;  1, -1,\;  1\!\!-\!k, -1,-1,\;  1,\;  k\!-\!\!1,\;  1,\;  k\!-\!\!1,\;  -1\big).
$$
for $k=2,3,\ldots$, and so $\mathbb Q$ certainly is one of the groups which from the perspective of Higman embeddings has a not very complicated set of  relations.

Then we suggest \textit{an algorithm} with some generic tools that allow to explicitly construct $\mathcal B$ by Higman's operations for groups from the listed classes, and more.

\medskip
For this algorithm in addition to the initial Higman operations we suggest a few extra \textit{auxiliary operations} which make our work with the Higman operations more convenient and intuitive, see Subsection 2.4 in \cite{The Higman operations and  embeddings}.

\medskip
We intentionally brought our example above for the rational group $\mathbb Q$ because this topic is related to Problem 14.10 (a) posted by Bridson and de la Harpe in Kourovka Notebook as a ``well-known problem''.

Namely they asked to explicitly embed the group $\mathbb Q$ of rational numbers into a finitely presented group. 

\medskip
Our algorithm in \cite{The Higman operations and  embeddings} gives the method how such an explicit embedding can be build, without displaying any explicit embedding though.
An explicit embedding was recently found in in \cite{Belk Hyde Matucci} by Belk, Hyde and Matucci.  \cite{Belk Hyde Matucci} also stresses that \cite{The Higman operations and  embeddings} \textit{``has described how to explicitly carry out Higman's construction for $\mathbb Q$ as well as many other groups of interest''.}

\medskip
In \cite{The Higman operations and  embeddings} we intentionally try to stay as close to Higman's technique with free constructions of groups as possible, and also to the methods of  \cite{A modified proof for Higman}. Some of the proof details could be shortened by using tricks with wreath products of groups from  \cite{Subnormal embedding theorems},
 \cite{metabelian},
 \cite{On abelian subgroups of},
\cite{The Criterion of Shmel'kin},  
 \cite{Subvariety structures}, 
 \cite{Mikaelian 2002}, \cite{Mikaelian 2003}, \cite{Mikaelian 2003b}, \cite{Mikaelian 2004}, \cite{Mikaelian 2005}, \cite{Mikaelian 2007} or \cite{Mikaelian 2017},
however, we refrain from mixing free constructions technique with wreath product methods.

\end{document}